The results for daily intakes of milk (in ounces) for ten randomly selected people were xˉ=18.23 and s=6.11. Find a 99% confidence interval for the population standard deviation σ. Assume that the distribution of ounces of milk intake daily is normally distributed.

What is the margin of error?
Answer #1

)solution

Given that,

= 18.23

s =6.11

n = 10

Degrees of freedom = df = n - 1 =10 - 1 = 9

At 99% confidence level the t is ,

= 1 - 99% = 1 - 0.99 = 0.01

/ 2 = 0.01 / 2 = 0.005

t /2  df = t0.005,9 = 3.250 ( using student t table)

Margin of error = E = t/2,df * (s /n)

= 3.250* ( 6.11/ 10) = 6.28